A person pushes with a force of 10.0 N on a stick
Answer #1

Here ,

1)

Pressure = force/area

Pressure = 10/(pi * (0.12 *10^-3)^2)

Pressure = 2.210 E+8 Pa

the pressure is A) 2.210 E+8 Pa

----------------------------------

2)

Here , for the apprent weight

apprent weight = Weight of object - weight of water displaced

apprent weight = 3000 - 0.25 * 1000 * 9.8

apprent weight = 550 N

the apprent weight is a) 550 N

3)

density = mass/volume

density = 28/(2.5 * 3.8 * 4) g/cm^3

density = 0.7368 g/cm^3

the density is c) 0.7368 g/cm^3

4.

as the density of plastic block is less than the density of water

it will a) Float